OpenStudy (anonymous):

The cost, C, in dollars, of fuel per month for Sanjay to operate his truck is given by C(v) = 0.0029v^2 - 0.48v +142, where v represents his average driving speed, in kilometres per hour. Find the most efficient speed at which Sanjay should drive his truck.

OpenStudy (anonymous):

First step that you have to do is to minimize the C in the equation. C(v)=0.0058v -0.48=0 v=0.48/0.0058 v=82.75862069 km/hr

OpenStudy (anonymous):

You cold also use a graph to find out that axis of sym is on 82.7 as well as the vertex.